3 less than 2 times a number is the same as 2 more than 3 times the number. Find the number
Answer by mananth(16949)

2x-3 = 3x+2
add +3 to both sides
2x-3+3 =3x+2+3
2x=3x+5add-3x
-3x+2x=3x-3x+5
-x= 5
divide by -1
x=-5